Solve for x. x - 1 3 0 4 x - 2 X = Submit Answer Tries 0/10
Answer #1

1:31 - - (-1)(x-2)-23 = 0 / 2X-X+2-12 -o 2 x^3 -10 = 0 - 5X +ZZ-10-0 => x(-5)+2(x)=a (x+(-5)=0 Z=-2 OF =5 R:-2.5